OpenStudy (anonymous):

cos(5/13-3/5) find the exact value

OpenStudy (anonymous):

\[\cos \left( \frac{ 5 }{ 13}-\frac{ 3 }{ 5 } \right)\]

OpenStudy (anonymous):

help

OpenStudy (campbell_st):

well you need 2 right triangles  next use the difference expansion for cos cos(X - Y) = cos(X)cos(Y) + sin(X)sin(Y) find sin(X) and sin(Y) from the triangles. hope it helps

OpenStudy (anonymous):

so the answer is 33/65

OpenStudy (campbell_st):

nope not the correct answer... notice how the expansion has a + sign....

OpenStudy (anonymous):

a= 12 but + sign I don't get

OpenStudy (campbell_st):

ok... wel its the difference equation for cos cos(A -B) = cos(A)cos(B) + sin(A)sin(B)

OpenStudy (anonymous):

cos(a)=5/13 cos(b)=-3/5+sin(a)=12/13 sin(b)=4/5 so than it is 18/65 if not how do you work it out